OpenStudy (icalibear):

Solve the system of equations using the substitution method –y + 3x = 6 y = –6x + 12

OpenStudy (anonymous):

might as well substitute y !!

OpenStudy (anonymous):

–y + 3x = 6 --> -y=6-3x --> y=-6+3x y = –6x + 12 _____________________________________________________ y =-6+3x y = –6x + 12

OpenStudy (anonymous):

–6x + 12 = -6+3x

OpenStudy (anonymous):

Now solve for X!!

OpenStudy (anonymous):